If z% of 60 is equal to 40% of 80, what is the value of z to the nearest tenth?
Found 2 solutions by jorel555, advanced_Learner:
Answer by jorel555(1290)

.4 x 80=32
60z=32
z=32/60=.53333=53.3333..% ☺☺☺☺

Answer by advanced_Learner(501)

.01z*60=0.4*80
0.6z=32
solve it